Calculate the pressure exerted by a 35 N cylinder with a base radius of 5 cm while resting on a floor
1
Expert's answer
2019-07-05T09:12:06-0400
"\rm{Pressure}=\frac{\rm{Force}}{\rm{Area}}"

"P=\frac{F}{A}=\frac{mg}{\pi r^2}"

"=\frac{35\:\rm{N}}{3.14\times (0.05\:\rm{m})^2}=4458\:\rm{Pa}"